John is 10 years older than Chris. In 2 years John will be twice as old as Chris. How old are they now?
Answer by Maths68(1474)

Let
Present age of John = j
Present age of Chris = c

John is 10 years older than Chris.
j=c+10 ..........(1)
In 2 years
John will be = j + 2
Chris will be = c + 2
Then,
John will be twice as old as Chris.
j + 2 = 2(c + 2)
j + 2 = 2c + 4
j = 2c + 4 - 2
J = 2c + 2 .............(2)
Put the value of j from (1) to (2)
c+10 = 2c + 2 .............(2)
c-2c=2-10
-c=-8
Multiply by -1 both sides
c=8
Put the value of c in (1)
j = 8 + 10
j = 18



Present age of John = j = 18 years old
Present age of Chris = c = 8 years old